OpenStudy (anonymous):

Mr. Wilson invested money in two accounts. His total investment was $9,000. If one account pays 4% in interest and the other pays 12% in interest, how much did he invest in each account if he earned a total of $680 in interest in 1 year? @iambatman

OpenStudy (kropot72):

Let the amount earning 4% be x. Let the amount earning 12% be y. x + y =9000 .............................(1) 0.04x + 0.12y = 680 ................(2) Now just solve the simultaneous equations. Can you do that?

OpenStudy (anonymous):

I think I can

OpenStudy (anonymous):

@iambatman batman i need you ! lol

OpenStudy (anonymous):

Look at @kropot72 explanation :P, it's the same thing as the last question just with more words.

OpenStudy (anonymous):

Exactly! It confuses me lol I don't even know which one to start with

OpenStudy (anonymous):

It doesn't matter, but you can pick equation one, since it's easier to isolate x or y.

OpenStudy (anonymous):

x=-y+9000?

OpenStudy (anonymous):

Yeah :D

OpenStudy (kropot72):

By rearranging (1) you can find y in terms of x as follows: y = 9000 - x Now substitute the value of y into equation (2) giving: 0.04x + 0.12(9000 - x) = 680 ...............(3) Can you multiply out equation (3) and combine like terms to solve for x?

OpenStudy (anonymous):

Oh goodness, I'm gonna have a heart attack over here lol

OpenStudy (anonymous):

You can do it either way cookii

OpenStudy (anonymous):

y=4000

OpenStudy (kropot72):

Correct!

OpenStudy (anonymous):

Now just sub that in one of the equations to get x :p

OpenStudy (anonymous):

x=5000!

OpenStudy (anonymous):

Pro

OpenStudy (anonymous):

whooooa, I've never done math so fast in my life! I have almost 10 minutes to finish this quiz!

OpenStudy (anonymous):

It's because I'm batman.

OpenStudy (anonymous):

lol is 12% 5000 & 4% is 4000?

OpenStudy (anonymous):

x = 4% y = 12% If you used Kropots explanation.